What is the amount of money that Lebo has to invest, at a simple interest rate of 12% per year,to earn R5 250,00 interest after three years?
1
Expert's answer
2021-10-06T13:33:47-0400

"interest=principal\times rate\times time\\525000=p\times\frac{12}{100}\times 3\\525000=0.36p\\p=1458333.3"